# Laminator: Verifiable ML Property Cards using Hardware-assisted Attestations

### v1

@@ III-B Desiderata for ML Property Attestations
    Efficient (low-overhead generation and verification of the attestation, even with large models and a wide range of useful ML properties).
@@
    Versatile (able to prove a wide variety of claims, and new claims created with minimal effort.)
Our goal is to design a framework to generate verifiable ML property cards using an attestation mechanism which can satisfy all the above requirements.
@@
Att 1 Attestation
where $a\overset{\$}{\leftarrow}\operatorname{Proc}()$ denotes probabilistic assignment (i.e. assigning to $a$ the value of a possibly-probabilistic procedure $\operatorname{Proc}()$).
@@
III-C Limitations of Prior Work
A popular ML property attestation mechanism is using cryptographic primitives such as zero knowledge proofs (ZKPs) and secure multiparty computation (MPC). However, none of the prior attestation scheme using cryptographic primitives satisfy all the three requirements.
  1. 1.
    ZKPs are scalable and can support multiple verifiers. However, they are computationally expensive and do not efficiently extend to neural networks. For instance, proof of training using ZKPs for a simple logistic regression model on a dataset of $2^{18}$ data records with 1024 attributes, and a batch size of 2014, takes 72 seconds for one epoch including training and proof generation time [17]. Finally, ZKPs are lack versatility, and require properties that can be adapted to the ZKP scheme [17], such as complex nonlinear activation functions.
  2. 2.
    MPC requires online interactions between the prover and verifier, meaning that it lacks scalability and efficiency, and requires training the model for each verification [11].
IV Verifiable ML Property Cards
ML property cards currently cover properties about datasets in the form of datasheets, and about models in the form of model cards. However, they currently do not include properties during each inference after deployment. To this end, we propose the notion of inference cards (Section IV-A). We then present different ML property attestations for different verifiable ML property cards (Section IV-BIV-C, and IV-D). Finally, we discuss how $\mathcal{V}$ can combine different attestations to make meaningful conclusions (Section IV-E).
IV-A Inference cards
Inference cards include properties for each inference: i) attesting that an output was from a specific model and input ii) verifiable inference for integrity of the outputs using ZKPs for ML models [40, 24, 16] and large language models [61] iii) attesting that model explanations, vector indicating the influence of different attributes in input to generate output, corresponds to a specific model, input, and output iv) attesting that an processed information was from either a pre-processing or post-processing filter for a specific input.
@@ IV-A Inference cards
Figure 1: Overview of ML property attestations and their relation to different verifiable model cards: attestations for datasheet (orange), model card (blue), inference card (red), along with external certificates (cyan).
IV-B Attestations for Datasheets
We first present dataset attestations required for datasheets: distributional property attestation of $\mathcal{D}_{tr}$. We hereafter refer to a hash function as $h(\cdot)$. All the frequently used notations are in Table I.
Distributional Property Attestation attests that distributional properties p (such as gender ratio) of $\mathcal{D}_{tr}$ by providing an assertion that it has some value which can later be checked with some requirements by $\mathcal{V}$ [11]. This can either be done during training [11, 17] or later during inference [11] with Algorithm 3. We can have different distributional properties across: i) a single attribute (e.g., male-to-female ratio) ii) multiple attributes (e.g., male-to-female and white-to-non-white ratios) iii) joint properties (i.e., fraction of data records for a given classification label for different subgroups). For a given $\mathcal{D}^{prov}_{tr}$, Laminator specifies the properties to be attested and generates the following: p, h($\mathcal{D}^{prov}_{tr}$), $sign_{sk}$(h($\mathcal{D}^{prov}_{tr}$), p)). We present the attestation in Algorithm 2.
@@ IV-B Attestations for Datasheets
Output: p, $\operatorname{Sign}_{\textrm{att}}(\texttt{'DistAtt'},\textsc{p},\texttt{h}(\mathcal{D}^{prov}_{tr}))$
@@
IV-C Attestations for Model Cards
We now cover different attestations for model cards which includes: i) proof of training which includes attesting $\mathcal{T}$ and $\mathcal{M}_{ar}$ ii) evaluation attestations (accuracy, on $\mathcal{D}^{prov}_{te}$ iii) on $\mathcal{D}^{prov}_{te}$ iv) robustness on $\mathcal{D}^{prov}_{te}$ We indicate these in blue (Figure 1).
Proof of Training attests that $\mathcal{M}$ was trained on $\mathcal{D}^{prov}_{tr}$. This requires training $\mathcal{M}$ using a training configuration $\mathcal{T}$ on $\mathcal{D}^{prov}_{tr}$ inside the trusted hardware. The attestation in shown in Algorithm 3.
@@ IV-C Attestations for Model Cards
Accuracy attestation indicates that $\mathcal{M}$ has a certain accuracy $\mathcal{A}_{acc}$ as measured using some dataset $\mathcal{D}^{prov}_{te}$. This ensures i) there are no false claims about $\mathcal{M}$’s $\mathcal{A}_{acc}$ ii) $\mathcal{M}$was evaluated on (a standardized) $\mathcal{D}^{prov}_{te}$ with no tampering. Laminator, in addition to $\mathcal{A}_{acc}$, generates $\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$) along with a signature binding them, i.e., $Sign_{\textrm{att}}$($\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$), $\mathcal{A}_{acc}$). We use an indicator function $\mathbb{I}_{\{a=b\}}$ that takes the value $1$ if $a=b$, and 00 otherwise. We summarize this in Algorithm 4.
@@
Output: $\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$), $\mathcal{A}_{acc}$, $\operatorname{Sign}_{\textrm{att}}(\texttt{'AccAtt'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{D}^{prov}_{te}),\mathcal{A}_{acc})$
@@
Fairness Attestation extends accuracy attestation to evaluate fairness of $\mathcal{M}$ on $\mathcal{D}^{prov}_{te}$ using different metrics such as accuracy parity, false positive and negative rates across subgroups. We refer to the fairness metric being evaluated as $\mathcal{A}_{fair}$. We attest that $\mathcal{M}$ satisfies $\mathcal{A}_{fair}$ on $\mathcal{D}^{prov}_{te}$ in Algorithm 5. We only indicate demographic parity but any other fairness metric can be used in its place.
@@
Output: $\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$), $\mathcal{A}_{fair}$,$\operatorname{Sign}_{\textrm{att}}(\texttt{'FairAtt'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{D}^{prov}_{te}),\mathcal{A}_{fair})$
@@
Robustness Attestation attests that $\mathcal{M}$ satisfies some acceptable level of robustness indicated by robust accuracy ($\mathcal{A}_{rob}$) on a test dataset containing adversarial examples ($\mathcal{D}_{rob}$). This proceeds in two steps: first, we generate $\mathcal{D}_{rob}$ along with its hash (h($\mathcal{D}_{rob}$)) inside the enclave. Since $\mathcal{D}_{rob}$ is generated locally, its hash acts as the certificate during verification. This certified that $\mathcal{D}_{rob}$ was created by adding adversarial examples generates with a perturbation of $\epsilon$ to $\mathcal{D}^{prov}_{te}$. In the second step, we use accuracy attestation (Algorithm 4) but using $\mathcal{D}_{rob}$ instead of $\mathcal{D}^{prov}_{te}$, yielding our robustness output.
@@
Output: $\mathcal{M}$), h($\mathcal{D}_{rob}$), $\mathcal{A}_{rob}$, $\operatorname{Sign}_{\textrm{att}}(\texttt{'RobustAttB'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{D}_{rob}),\mathcal{A}_{rob})$
@@
IV-D Attestations for Inference Cards
@@
Input-Model-Output Attestation (IOAtt), shown in Algorithm 7 attests that the output $\mathcal{O}$ was generated from $\mathcal{M}$ for a given input $\mathcal{I}$. This ensures i) integrity against malicious $\mathcal{P}$ who can tamper with $\mathcal{O}$ ii) non-repudiation to prevent a malicious client from claiming that $\mathcal{O}$ does not belong to $\mathcal{I}$ iii) false claim that $\mathcal{O}^{\prime}$ is from $\mathcal{M}$ for $\mathcal{I}$ instead of $\mathcal{O}$.
@@
Output: $\mathcal{O}$, $\mathcal{M}$), $\mathcal{I}$), h($\mathcal{O}$), $\operatorname{Sign}_{\textrm{att}}(\texttt{'IOAtt'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{I}),\texttt{h}(\mathcal{O}))$
@@
IV-E Chain of Attestations and External Certificates
In addition to the above attestations, there could be additional certificates which allow verifiers to check the integrity of datasets and trained models. We consider two different external certificates: i) training dataset ii) test dataset, which certify the quality of these dataset as indicated by some trusted organization (e.g., CIFAR). Combining these external certificates can help $\mathcal{V}$ conclude that the model was trained and evaluated on dataset considered good for a particular task.
The assertions of different attestations and external certificates by themselves might not be useful in practice. For instance, input-model-output attestation suggests that for an input $\mathcal{I}$, $\mathcal{M}$ generates the output $\mathcal{O}$. However, this information is useful only if $\mathcal{V}$ knows how good $\mathcal{M}$ is in terms of accuracy, fairness or robustness. In order to use attestations in practice, we chain together different attestations included to obtain verifiable ML property cards that allow us to arrive at a wider range of conclusions.
As a motivating example, consider a company which has outsourced training to a third party before deploying some model ($\mathcal{M}$) as a service to clients. Using input-model-output attestation, the company can indicate that for their input $\mathcal{I}$, $\mathcal{M}$ generated $\mathcal{O}$ (Algorithm 7). Additionally, they have to attest that $\mathcal{M}$ is a good model in terms of having “good” accuracy on on some $\mathcal{D}_{te}$ using accuracy attestation (Algorithm 4). This helps the clients or $\mathcal{V}$ conclude that their $\mathcal{O}$ is indeed from a good model. Additionally, including other attestations such as fairness (Algorithm 5) and robustness can help show that $\mathcal{M}$ satisfies other properties as required by different applications. In all these cases, including the certification for $\mathcal{D}_{te}$ will also help conclude that the metrics are computed on a standardised test dataset which is certified by a trusted organization.
@@ IV-E Chain of Attestations and External Certificates
V Laminator: Framework Design
@@
We present Laminator, a framework to generate verifiable ML property cards using ML property attestation. We identify hardware-assisted ML property attestations using Trusted Execution Environments (TEEs) as the choice for realizing efficient, scalable and versatile ML property attestation. TEEs already have the notion of hardware-assisted remote attestation to prove local system or software configuration to a remote verifier. The trusted computing research community has extended this to the notion of property attestation [54], and recent developments by hardware vendors [26, 49], make it possible to run and train ML models efficiently inside TEEs. We are the first to propose a software framework for hardware-assisted ML property attestation.
@@
Laminator establishes bindings between key components: the model and its inputs/outputs, the model and its accuracy with respect to a test dataset, the model and its training dataset and configuration, and distributional properties of a dataset. Bindings are signed with the TEE’s secret attestation key, yielding attestations. As these can be generated by anyone with the requisite hardware, and validated by anyone, this approach is inherently scalable. They enable verifiers to draw conclusions about the model and training dataset properties during training and inference. We implement the different attestations described in Section IV. We revisit the completeness and how to extend Laminator for new attestations in Section VIII. We now discuss the design and implementation of Laminator for $\mathcal{P}$ and $\mathcal{V}$.
SGX EnclavesMeasurerPytorchPythonGramine LibOSAttestationsEndorserEndorsementsAssertionBundleIntel APICertificatesVerifierDCAPProperty Cards
@@
Figure 2: Overview of Laminator’s design. Components already existing or part of the infrastructure are indicated in gray while components implemented are indicated in orange. There can be different enclaves for different attestations which are generated by changing the “measurer” (in green). Dashed lines correspond to components for endorser while solid lines correspond to components from TEE.
@@
We present an overview of Laminator’s implementation in Figure 2. An SGX enclave runs a Python script (“measurer”) that measures properties of a model, dataset, or inference, and produces an attestation of a property card fragment (i.e. a chunk of JSON with the same structure as model card metadata that will be combined with other such fragments to produce the full property card). A combination of external certificates and various enclaves’ attestations is referred to as the assertion bundle. Since it is possible to endorse new enclaves without updating all verifiers’ software, this approach is inherently flexible. Furthermore, provisioning certificates are included in the enclave platforms using an Intel API to later check for validity of the attestations. We also consider assertions by endorsers (e.g., CIFAR, Huggingface) of the existing kind, who can be trusted to provide external certificates asserting that some publicly available models or datasets are of good quality, or simply to name a dataset. During verification, the assertion bundle can be passed to $\mathcal{V}$ who verifies the attestations using certificates from the TEE manufacturer (in this case, obtained from an API provide by Intel and used by the DCAP Quote Verification Library), and combines their assertions with those of the endorsers to obtain a set of verified ML property cards.
V-A The Prover

### v3

@@ 3.2. Desiderata for ML Property Attestations
    Efficient (low-overhead generation of attestations).
@@
    Versatile (supports a wide variety of properties; requires minimal effort to add support for new properties).
Our goal is to design a framework to generate verifiable ML property cards that can satisfy all the above requirements.
@@
In the rest of the paper, we will describe attestations in the form shown in Algorithm 1.
@@
where $a\overset{\$}{\leftarrow}\operatorname{Proc}()$ denotes probabilistic assignment (i.e. assigning to $a$ the value of a possibly-probabilistic procedure $\operatorname{Proc}()$). Not all attestations necessarily include all fields.
Attestation 1 Sample attestation
@@
3.3. Limitations of Prior Work
ML property cards can be made verifiable using cryptographic techniques like zero-knowledge proofs (ZKPs) or secure multiparty computation (MPC) for verifiable training (Sun and Zhang, 2023; Shamsabadi et al., 2023; Garg et al., 2023; Abbaszadeh et al., 2024; Eisenhofer et al., 2022) and inference (Liu et al., 2021; Feng et al., 2021; Weng et al., 2021; Huang et al., 2022; Zhang et al., 2023; Ghodsi et al., 2017). But no prior scheme meets all three requirements.
  1. (1)
    ZKPs are scalable and can support multiple verifiers. However, they are computationally expensive and do not efficiently extend to neural networks. For instance, proof of training using ZKPs for a simple logistic regression model on a dataset of $2^{18}$ data records with 1024 attributes, and a batch size of 2014, takes 72 seconds for one epoch including training and proof generation time (Garg et al., 2023). Finally, ZKPs lack versatility, as they only support properties that can be adapted to the ZKP scheme (Garg et al., 2023), limiting their use for complex functions (e.g., nonlinear activation).
  2. (2)
    MPC requires online interactions between the prover and verifier, meaning that it lacks scalability and efficiency, and requires retraining the model for each verification (Duddu et al., 2024).
Despite ongoing advances in cryptographic techniques (Ng and Chow, 2023), purely cryptographic protection of many ML operations (e.g., training), remain inefficient. Recent optimizations for TEEs to run ML workloads efficiently (Nvidia, 2024; Zhu et al., 2024; Intel, 2024) as well as secure outsourcing protocols to GPUs (e.g., (Ng et al., 2021; Mo et al., 2022)) offer efficient solutions by combining TEEs with hardware accelerators. In light of these, we explore the use of TEE remote attestation for ML properties, with negligible overhead compared to a baseline setting where an ML application is already running in a confidential computing setting. We conjecture that using TEE attestation, instead of mechanisms like ZKP, makes verification more efficient. We discuss other TEE-based work which focus on confidentiality instead of integrity in Section 8.
4. Verifiable ML Property Cards
Current ML property cards describe properties about datasets in the form of datasheets, and about models in the form of model cards. However, they currently do not include properties during each inference after deployment. To this end, we propose the notion of inference cards (Section 4.1). We then present different ML property attestations for different verifiable ML property cards (Section 4.24.3, and 4.4) by carrying out operations inside a TEE and then including an attestation in the ML property card. Finally, we discuss how $\mathcal{V}$ can combine different attestations to make meaningful conclusions (Section 4.5).
4.1. Inference cards
Inference cards include properties for each inference, for instance, showing that an output was obtained from a specific model with a particular input and, when using an explainable model, that an explanation derives from the same inference (Guidotti et al., 2018). Prior works using non-interactive ZKPs for verifiable inference in ML models (Liu et al., 2021; Huang et al., 2022; Feng et al., 2021) and large language models (Sun et al., 2024), aim to provide similar functionality to inference cards.
@@ 4.1. Inference cards
Figure 1. Overview of ML property attestations and their relation to different verifiable model cards: attestations for datasheet (orange), model card (blue), inference card (red), along with external certificates (cyan).
4.2. Attestations for Datasheets
We first present dataset attestations required for datasheets: distributional property attestation of $\mathcal{D}_{tr}$. We hereafter refer to a hash function as $h(\cdot)$. All the frequently used notations are in Table 1.
Distributional Property Attestation attests that distributional properties p (such as gender ratio) of $\mathcal{D}_{tr}$ by providing an assertion that it has some value which can later be checked with some requirements by $\mathcal{V}$ (Duddu et al., 2024). This can either be done during training (Duddu et al., 2024; Garg et al., 2023) or later during inference in combination with Algorithm 3 (Duddu et al., 2024). We can have different distributional properties across: i) a single attribute (e.g., male-to-female ratio) ii) multiple attributes (e.g., male-to-female and caucasian-to-non-caucasian ratios) iii) joint properties (i.e., fraction of data records for a given classification label for different subgroups). For a given $\mathcal{D}^{prov}_{tr}$, Laminator specifies the properties to be attested and generates the following: p, h($\mathcal{D}^{prov}_{tr}$), and $sign_{sk}$(h($\mathcal{D}^{prov}_{tr}$), p)). We present this in Algorithm 2.
@@ 4.2. Attestations for Datasheets
Output: p, $\operatorname{Sign}_{\textrm{att}}(\texttt{`DistAtt'},\textsc{p},\texttt{h}(\mathcal{D}^{prov}_{tr}))$
@@
4.3. Attestations for Model Cards
We now cover different attestations for model cards, including: i) proof of training, which includes attesting $\mathcal{T}$ and $\mathcal{M}_{ar}$ ii) evaluation attestations, which includes accuracy, fairness, and robustness on $\mathcal{D}^{prov}_{te}$. We indicate these in blue (Figure 1).
Proof of Training attests that $\mathcal{M}$ was trained on $\mathcal{D}^{prov}_{tr}$. This requires training $\mathcal{M}$ using a training configuration $\mathcal{T}$ on $\mathcal{D}^{prov}_{tr}$. The attestation in shown in Algorithm 3.
@@ 4.3. Attestations for Model Cards
Accuracy attestation indicates that $\mathcal{M}$ has a certain accuracy $\mathcal{A}_{acc}$ as measured using some dataset $\mathcal{D}^{prov}_{te}$. This ensures i) there are no false claims about $\mathcal{M}$’s $\mathcal{A}_{acc}$, and ii) $\mathcal{M}$was evaluated on (a standardized) $\mathcal{D}^{prov}_{te}$ with no tampering. Laminator, in addition to $\mathcal{A}_{acc}$, generates $\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$) along with a signature binding them, i.e., $Sign_{\textrm{att}}$($\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$), $\mathcal{A}_{acc}$). We use an indicator function $\mathbb{I}_{\{a=b\}}$ that takes the value $1$ if $a=b$, and 00 otherwise. We summarize this in Algorithm 4.
@@
Output: $\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$), $\mathcal{A}_{acc}$, $\operatorname{Sign}_{\textrm{att}}(\texttt{`AccAtt'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{D}^{prov}_{te}),\mathcal{A}_{acc})$
@@
Fairness Attestation extends accuracy attestation to evaluate fairness of $\mathcal{M}$ on $\mathcal{D}^{prov}_{te}$ using different metrics such as accuracy parity, false positive and negative rates across subgroups. We refer to the fairness metric being evaluated as $\mathcal{A}_{fair}$. We attest that $\mathcal{M}$ satisfies $\mathcal{A}_{fair}$ on $\mathcal{D}^{prov}_{te}$ in Algorithm 5. We use demographic parity as an example, but any fairness metric can be substituted.
@@
Output: $\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$), $\mathcal{A}_{fair}$,$\operatorname{Sign}_{\textrm{att}}(\texttt{`FairAtt'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{D}^{prov}_{te}),\mathcal{A}_{fair})$
@@
Robustness Attestation attests that $\mathcal{M}$ satisfies some acceptable level of robustness indicated by robust accuracy metric ($\mathcal{A}_{rob}$) on a test dataset containing adversarial examples ($\mathcal{D}_{rob}$). This proceeds in two steps: first, we generate $\mathcal{D}_{rob}$ along with its hash (h($\mathcal{D}_{rob}$)) inside the enclave. Since $\mathcal{D}_{rob}$ is generated locally, its hash serves as the certificate during verification. This certifies that $\mathcal{D}_{rob}$ was created by adding adversarial examples, generated with a perturbation of $\epsilon$, to $\mathcal{D}^{prov}_{te}$. In the second step, we use accuracy attestation (Algorithm 4) but using $\mathcal{D}_{rob}$ instead of $\mathcal{D}^{prov}_{te}$, yielding $\mathcal{A}_{rob}$.
@@
Output: $\mathcal{M}$), h($\mathcal{D}_{rob}$), $\mathcal{A}_{rob}$, $\operatorname{Sign}_{\textrm{att}}(\texttt{`RobustAttB'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{D}_{rob}),\mathcal{A}_{rob})$
@@
4.4. Attestations for Inference Cards
@@
Input-Model-Output Attestation (IOAtt), shown in Algorithm 7 attests that the output $\mathcal{O}$ was generated from $\mathcal{M}$ for a given input $\mathcal{I}$. This ensures i) integrity against malicious $\mathcal{P}$ who can tamper with $\mathcal{O}$ ii) non-repudiation to prevent $\mathcal{P}$ from claiming that $\mathcal{O}$ is not from to $\mathcal{I}$ iii) false claim that $\mathcal{O}^{\prime}$ is from $\mathcal{M}$ for $\mathcal{I}$ instead of $\mathcal{O}$.
@@
Output: $\mathcal{O}$, $\mathcal{M}$), $\mathcal{I}$), h($\mathcal{O}$), $\operatorname{Sign}_{\textrm{att}}(\texttt{`IOAtt'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{I}),\texttt{h}(\mathcal{O}))$
@@
4.5. Chains of Attestations and Certificates
Along with the above attestations, certificates from trusted certifiers (e.g., CIFAR) can help verifiers to validate the quality of the training or test datasets. These certificates can certify a non-computational property such as e.g. data provenance or suitability for a particular application. Combining these certificates with attestations can help $\mathcal{V}$ conclude that the model was trained and evaluated on dataset considered good for a particular task.
The assertions of different attestations along with the external certificates by themselves might not be useful in practice. For instance, input-model-output attestation suggests that for an input $\mathcal{I}$, $\mathcal{M}$ generates the output $\mathcal{O}$. However, this information is useful only if $\mathcal{V}$ knows how good $\mathcal{M}$ is in terms of accuracy, fairness or robustness. To use attestations in practice, we chain multiple attestations, used to create verifiable ML property cards, allowing us to arrive at a broader range of conclusions.
As a motivating example, consider a company testing $\mathcal{M}$ for “good” accuracy on $\mathcal{D}_{te}$ using accuracy attestation (Algorithm 4). This helps $\mathcal{V}$ conclude that their $\mathcal{O}$ is indeed from a good model. Including additional attestations, such as fairness (Algorithm 5) and robustness (Algorithm 6), demonstrates that $\mathcal{M}$ meets other application-specific requirements. In all cases, certifying $\mathcal{D}_{te}$ ensures that metrics are computed on a standardized test dataset certified by a trusted organization.
@@ 4.5. Chains of Attestations and Certificates
5. Laminator: Framework Design
@@
We present Laminator, a framework to generate verifiable ML property cards using ML property attestation from TEEs. TEEs already have the notion of hardware-assisted remote attestation to prove local system or software configuration to a remote verifier. The trusted computing research community has extended this to the notion of property attestation (Sadeghi and Stüble, 2004), and recent developments by hardware vendors (Intel, 2024; Nvidia, 2024), make it possible to train and run ML models efficiently inside TEEs. We are the first to propose a software framework for hardware-assisted ML property attestation.
@@
Laminator establishes bindings between key components: (i) the model and its inputs/outputs (ii) the model and its accuracy with respect to a test dataset (iii) the model and its training dataset and configuration (iv) and distributional properties of a dataset. Bindings are signed with the TEE’s secret attestation key, producing attestations. Since these can be generated and validated by anyone with the requisite hardware, this approach is inherently scalable and allows verifiers to draw conclusions about the model and training dataset properties. We implement the different attestations described in Section 4. We revisit the completeness and how to extend Laminator for new attestations in Section 8. We now discuss the design and implementation of Laminator for $\mathcal{P}$ and $\mathcal{V}$.
Trusted CertifierSEV-SNP Virtual MachineSGX EnclaveMeasurerPytorchPythonGramine LibOSAttestationsTrusted CertifierNon-computationalProperty CertificatesAssertionBundleVerifierProperty Cards
@@
Figure 2. Overview of Laminator’s design: Components already existing or part of the infrastructure are indicated in gray while components specific to Laminator are indicated in orange. There can be different enclaves for different attestations which are generated by changing the “measurer” (in green). Dashed lines correspond to assertions by a trusted certifier, while solid lines correspond to assertions by a TEE.
@@
We present an overview of Laminator’s implementation in Figure 2. A TEE runs a Python script (“measurer”) that measures properties of a model, dataset, or inference, and produces an attestation of a property card fragment (i.e. a chunk of JSON with the same structure as model card metadata that will be combined with other such fragments to produce the full property card). The combination of certificates and attestations is called the assertion bundle. This approach can be made highly versatile, as new enclaves can be endorsed without updating all verifiers’ software.
We also consider assertions by trusted certifiers (e.g., CIFAR, Huggingface) who do not take advantage of remote attestation, but who can be trusted to provide external certificates asserting that some publicly available models or datasets are of good quality, or simply to name a dataset. During verification, the assertion bundle can be passed to $\mathcal{V}$ who verifies the attestations using certificates from the TEE manufacturer, and combines them to obtain a set of verified ML property cards.
We implement our proof of concept using Intel’s SGX as a TEE platform, but the design is not dependent on platform-specific details. It can be equally implemented on any other TEE with remote attestation support. Different implementations can interoperate as long as the verifier can validate each TEE’s remote attestations.
5.1. The Prover